OpenStudy (anonymous):

The question is: The radius of a circular disk is given as 22 cm with a maximum error in measurement of 0.4 cm. Use differentials to estimate the maximum possible error in the calculated area of the disk. I have determined the area to be 55.292cm^2 however when I input the value to calculate the percent error, I am getting an incorrect value. I have tried dA/A = 55.292/22cm but its not giving me the right value. Anyone help?

OpenStudy (anonymous):

is there a diagram on how the circular disk looks like?

OpenStudy (anonymous):

the volume of a circle is V=pi*r^2 therefore we would have to use that formula. I have already determined the change in the area, the only thing is getting the percent error.

OpenStudy (aum):

\[ A = \pi r^2 \\ \frac{dA}{dr} = 2\pi r \\ \frac{\Delta A}{\Delta r} \approx 2\pi r \\ \Delta A \approx 2\pi r * \Delta r \\ r = 22; ~~~\Delta r = 0.4 \\ \Delta A \approx 2\pi * 22 * 0.4 = 17.6\pi = 55.2920 \text{ cm}^2. \\ \]

OpenStudy (aum):

Do they ask for the percent error or the maximum possible error? The maximum possible error = 55.2920 cm^2 rounded to one decimal of 55.3 cm^2.

OpenStudy (anonymous):

asking for Maximum error.

OpenStudy (aum):

55.3

OpenStudy (aum):

Maximum possible error = 55.3 cm^2 Area of the circle = pi * r^2 = pi * (22)^2 = 1520.5 cm^2 Percent error = 55.3/1520.5 * 100 = 3.64% But if they did not ask for percent error then don't use this answer.

OpenStudy (anonymous):

Thank you for your help!